### v1

@@ 4 Results
As outlined in section LABEL:label:Framework, we evaluate conventional deep learning models using performance metrics which take into account the catchments, prediction horizon, and window length of the input time series. We evaluate four distinct catchment modelling strategies using a conventional LSTM model to identify the optimal catchment modelling strategy using a subset of the data from South Australia as a case study. We compare the Individual, Batch-Indicator, Batch-Static and Ensemble catchment modelling strategies. The analysis considers metrics such as SER(with percentile cuts at 1, 2, 5, 10, 50, and 75) and RMSE mean across five prediction horizons. This enables a thorough evaluation the influence of extreme flood conditions while making comparisons across different models and architectures. Table 3 presents the results, highlighting that the Individual catchment strategy emerged as the most effective across various metrics (SER percentiles). This strategy has the lowest (best) RMSE, which indicates that separate models for individual gauging stations gives the best performance. This is because each catchment has unique characteristics that could be based on weather patterns, climate, and topography. In Table 3, the Stacked Ensemble model gives poor performance, which is not even able to beat the individual LSTM models. This could be due to problems in combining the data from the Temporal with the Static model.
@@
The next step is to compare different deep learning models for catchments in South Australia, including the LSTM, BD-LSTM, CNN, ED-LSTM, and Quantile-LSTM. We chose South Australia dataset for comprehensive evaluation since it has a lower number of gauging stations and data when compared to large states such as New South Wales. As shown in Table 4, we find that the Quantile-LSTM provides the best prediction accuracy (RMSE of 0.0929) and SER percentiles. This is close to the LSTM model (RMSE of 0.0963), and CNN (RMSE of 0.0976), followed by BD-LSTM and ED-LSTM which give the lowest performance, respectively.
@@
Based on the analysis in South Australia, the Quantile-LSTM provides the best results and thus we proceed to employ it across all states in Australia. Table 5 presents the results, highlighting that in some states, the prediction accuracy (RMSE) is better than others, which is correlated to the SER percentiles. The difference in the results across the states could be due to different levels of extreme values in the data, and also the size of the data and the presence of noise. The Northern Territory and Western Australia have the worst model accuracy (RMSE) while New South Wales, Australian Capital Territory, and Queensland, and Victoria have some of the best prediction accuracies (Table 5). Notably, New South Wales and Queensland feature the highest number of catchments (gauging stations) and are states that are prone to extreme floods.
Finally, in Figures 12 to 16, we provide a visualisation of actual and predicted values using Quantile-LSTM for selected gauging stations (using Basin ID) for the different states for a period of more than two years. In most of the figures, we notice that short duration extreme rainfall events that contribute to high flood probabilities are predicted well by the Quantile-LSTM model. However, in most cases (eg. Figure 9 2010-07 - 2010-10 and Figure 12 2012-04 - 2012-07), the Quantile-LSTM model provides a lower flood probability when the actual flood probability is very large. This is noticeable in most of the selected Basin IDs for the respective states. This bias in prediction can potentially be corrected with a weighting mechanism, data augmentation or regularisation mechanism in the deep learning models.
Figure 9: South Australia Basin ID: A5030502
Figure 10: New South Wales Basin ID: 401009
Figure 11: Queensland Basin ID: 919309A
Figure 12: Victoria Basin ID: 401203
Figure 13: Australian Capital Territory Basin ID: 410730
Figure 14: Western Australia Basin ID: 604053
Figure 15: Northern Territory Basin ID: G8200045
Figure 16: Tasmania Basin ID: 318076
@@ 5 Discussion
In the first set of experiments, we compared the performance of the modelling strategies at the catchment scale using a conventional LSTM model for five step (day)-ahead prediction of streamflow (flood probability) using climate data. We found that the Individual catchment modelling strategy, where a separate model is used for each gauging station data, provided the best results (Table 3). This is because each catchment has unique characteristics that could be based on weather patterns, climate, and topography. In Table 3, the Stacked Ensemble model gives poor performance, which is not even able to beat the individual LSTM models. This could be due to problems in combining the data from the Temporal with the Static model. Generally, it is difficult for a single machine learning model to represent a wide range of data, even if they may be from the same domain, such as in a regionalization context. Regionalization would be a challenge for machine learning in general, as a single model cannot be used (fitted to) for several tasks, and in this case, the local catchments can be seen as different tasks as they can have differences in the data distribution dependent on climate/weather/geography. Approaches such as multitask and transfer learning [86, 87] could be considered, where some underlying model knowledge can be re-used and applied to related datasets, i.e. catchments in this case.
In the case of deep learning models using the Individual catchment strategy (Table 4), we observe that the Quantile-LSTM model provides the best results. However, Conventional LSTM and CNN models provide accuracy close to Quantile-LSTM model, which specifically includes a mechanism for quantifying uncertainty. Although designed for image processing purposes, the CNN reproduces streamflow better than the LSTM models that were designed for temporal sequences (it captures windows well that have little streamflow, as well as windows that contain extreme streamflows). CNNs may also have advantages for capturing spatial regions (regionalisation) due to their capabilities in image processing, but LSTM models were the best due to their capacity to simulate temporal data. The performance of these types of models can however be comparable as was shown in previous studies, especially for multi-step ahead prediction [56]. Furthermore, the Individual catchment strategy does not incorporate the spatial catchment data into the deep learning models, although multivariate data of the local catchment was used. Hence, this could be another reason why LSTM models outperformed the CNN. ED-LSTM and BD-LSTM, in this case, did not provide the a good performance, as was highlighted in past studies for univariate multi-step ahead prediction [56]. This could be due to the multivariate nature of the data that features extremes.
The results, reinforce the efficacy of the Quantile-LSTM as a robust and adaptable choice for hydrological modelling, specifically for South Australia. This performance highlights the Quantile-LSTM model resilience and effectiveness in capturing both extreme and regular flood occurrences, thereby reinforcing its suitability for capturing long-term hydrological patterns. However, as indicated, the Quantile-LSTM model can be biased in the extremes. This bias in prediction can potentially be corrected with a weighting mechanism, data augmentation or regularisation mechanism in the deep learning models.
The prediction of the future steps essentially gives an outlook - up to five days ahead, which would be useful for extreme events. However, the performance accuracy deteriorates as the prediction horizon increases in multi-step ahead problems (RMSE, SER). The results deteriorate as the gap in the values to predict increases with increasing prediction horizon.
@@
In our initial experiments (Table 3), we found the Individual catchment strategy to have the best performance. Interestingly, the Batch-Static strategy uses the same static features (7 summary features, stated in Table 1 ) as the Ensemble strategy which performs much worse (RMSE and SER percentiles). It seems that the Stacked Ensemble strategy cannot meaningfully extract and combine the static and temporal information from the respective models (Table 3). Comparatively, The Batch-Indicator Model is essentially the Batch Model without the inclusion of static features. It operates on the same time-series features as the Individual Architecture, but with a collective approach, running a single model for all stations in the state rather than individual models for each station. This consolidation aims to optimize computational resources.. The Stacked Ensemble strategy first generated activations from the temporal data using one model, and then generated activations from the static data in another model to more accurately capture temporal and static information in an ensemble. The Individual strategy features a specific fit to one catchment at a time and is unaffected by information from other catchments. For example, catchments might vary due to variations in weather patterns, climate, and topography. Overall, these findings are consistent with the work of Leontjeva et al. [24] who similarly compared a variety of strategies i) standalone model using univariate temporal data ii) standalone model using bimodal data (temporal and static) iii) ensemble model using bimodal data. In that study the standalone model using bimodal data fails to capture information from both sources, while the performance of the ensemble model depended on the distribution of the data.
@@
Despite the success of evaluating several machine learning approaches for predicting flood extremes, this study has some obvious limitations. Firstly, the models do not naturally project model uncertainties into the predictions, which is vital for natural disasters caused by extreme events. This can be addressed by Bayesian deep learning [88, 89]. Secondly, we have not accounted for topography information in the data to develop a robust spatiotemporal model.
@@
In the case of larger catchments, the steepness and variation in topography can influence rainfall patterns, and particularly the within catchment distribution of the rainfall. In addition, topography and catchment shape can influence the timing between rainfall and flood extremes. Implicitly, the LSTM models should be able to capture these variations for the individual catchment strategy based on the data, but it is possible that adding additional information will assist the model in identifying the patterns. We have also did not incorporate the changes in climate conditions due to El Nino climate effect, which is a longer scale pattern in the Australian rainfall [90, 91]. Again, the LSTM should be able to discover this from the data, given a long enough time series. However, adding this as a variable might assist the model identifying the variations in the data. Finally, the overall data set does not have the same length and missing data across the individual stations. The data is also sparse in space given the locations of the gauging stations (catchments), and this means if we need data for specific regions where no gauging stations are available, generative machine learning models [92], and general circulation models [93] could be used.
@@ 6 Conclusions
In this comprehensive study, we undertook a critical evaluation of existing deep learning models and introduced innovative methodologies for the accurate prediction of streamflow extremes across numerous catchments. Leveraging a multi-step-ahead forecasting strategy, our investigation revealed that a tailored approach, where each catchment is paired with a bespoke model dedicated to its distinct gauging station, significantly elevates the accuracy of extreme value forecasts. It’s important to highlight that this precision in modeling ensures that local variability and unique hydrological characteristics are captured effectively, providing a more nuanced understanding of each catchment’s behavior in the face of extreme weather events.
Moving to the forefront of our findings, we unveiled the potential of our conditional ensemble model, named Quantile-LSTM, which integrates quantile regression to surpass all baseline machine learning models in terms of predictive error margins for streamflows. This model emerges as a superior tool not only due to its performance but also because it represents a leap towards a more reliable prediction by embracing the inherent uncertainties in hydrological forecasting. Looking ahead, we recognize that the path to impeccable forecasts is iterative and progressive. Our research underscores the vast room for improvement that can be addressed by extending our dataset and refining our model, particularly in the area of uncertainty quantification. The aspiration is to reach a point where our model can offer robust predictions with quantified confidence levels, transforming how stakeholders prepare for and respond to the challenges posed by climate-induced flood extremes.
@@
    Developers: Siddharth Khedkar and Rohitash Chandra
@@
S. Khedkar and J. Ng contributed to coding, writing and experiments. R. Vervoort contributed to conceptualisation, writing and supervision. R. Chandra contributed to conceptualisation, writing, code analysis, and supervision.

### v2

@@ 4 Results
We first evaluate four distinct catchment modelling strategies using a conventional LSTM model to identify the optimal catchment modelling strategy using a subset of the data from South Australia as a case study. We compare the Individual, Batch-Temporal, Batch-Static and Ensemble catchment modelling strategies. The analysis considers metrics such as SER(with percentile cuts at 1, 2, 5, 10, 50, and 75) and RMSE mean across five prediction horizons. This enables a thorough evaluation of the influence of extreme flood conditions while making comparisons across different models and architectures. Table 3 presents the results, highlighting that the Individual catchment strategy emerged as the most effective across various metrics (SER percentiles). This strategy has the lowest (best) RMSE, which indicates that separate models for individual gauging stations provide the best performance.
@@
The next step is to compare deep learning models for catchments in South Australia, including the LSTM, BD-LSTM, CNN, ED-LSTM, and Ensemble Quantile-LSTM. We chose the South Australia dataset for comprehensive evaluation since it has a lower number of gauging stations and data when compared to large states such as New South Wales. As shown in Table 4, we find that the Quantile-LSTM (q=0.05) provides the best prediction accuracy (RMSE of 0.0801) and SER with high $\alpha$ values (50 and 75). This is close to the LSTM model (RMSE of 0.0963), and CNN (RMSE of 0.0976), followed by BD-LSTM and ED-LSTM which give the lowest performance, respectively. The Quantile-LSTM (q=0.95) which aims to capture the upper quantile of the predictive interval performs the best near the peaks (SER 1% and SER2%). The results reinforce the efficacy of the Ensemble Quantile-LSTM model as a robust choice for hydrological modelling with uncertainty quantification in South Australia. This performance highlights the resilience and effectiveness in capturing extreme and regular streamflow trends, where extreme trends could be associated with flood occurrences.
@@
Furthermore, we extend the Ensemble Quantile-LSTM model across all states in Australia. We only evaluate the performance of Ensemble Quantile-LSTM models on 5 selected basins from each state; i.e. we selected the stations with the highest runoff ratio in each state. Table 5 presents the results, highlighting that in some states, the prediction accuracy (RMSE) is better than others, which is correlated to the SER percentiles. The difference in the results across the states could be due to different levels of extreme values in the data, and also the size of the data and the presence of noise. The Northern Territory (NT) and Western Australia (WA) have the worst model accuracy (RMSE), while NSW, Australian Capital Territory (ACT), Queensland (QLD), and Victoria (VIC) have some of the best prediction accuracies (Table 5). Notably, NSW and Queensland feature the highest number of catchments (gauging stations) and are also the states that are prone to extreme floods [90].
In Figures 6 to 7, we provide a visualisation of actual and predicted values using the Ensemble Quantile-LSTM model for selected gauging stations (showing Basin ID) for the different states for one year. In most cases, we notice that short-duration extreme rainfall events that contribute to a high chance of flooding are well captured by the confidence interval predicted using the Ensemble Quantile-LSTM model. However, there are some cases where the actual streamflow crosses the 90% confidence interval predicted by the Ensemble Quantile-LSTM model (e.g. Figure 6(a) and 6(c)). We also notice that in some basins where the flow is mostly zeros, the Quantile-LSTM (q=0.05) is biased towards zero flows (e.g. Figure 6(c) and 7(c)).
Finally, Figure 8 and 9 present the 5-day average of actual and predicted streamflow and associated flood risk indicators for 4 selected catchments in New South Wales, Queensland, South Australia, and Western Australia, respectively, using the Ensemble Quantile-LSTM model. The first catchment (8(b)), Barambah Creek River at Litzows, exhibits multiple flood events over the study period, with both actual and predicted streamflows exceeding the flood threshold on multiple occasions. The predicted streamflow uncertainty, represented by the 90% confidence interval, captures most observed peaks, indicating the model’s ability to anticipate extreme events. The associated flood risk indicator can capture the two major actual flooding events in early 2011 and 2013, respectively. Similarly, for Nepean River at Maguires Crossing, the model successfully identifies 3 of the 4 major flood events, albeit with some variability in prediction accuracy. Overall, these results demonstrate the model’s capability to capture flood dynamics across diverse hydrological conditions, which is critical for early warning and disaster management. A similar performance is observed with Cullyamurra Water Hole (Figure 9(a)) and Darkin River at Pine Plantation (Figure 9(b)).
(a) South Australia Basin ID: A0030501
(b) New South Wales Basin ID: 401009
(c) Queensland Basin ID: 919309A
(d) Victoria Basin ID: 401203
Figure 6: Ensemble Quantile-LSTM streamflow predictions for a randomly selected basins from South Australia, Queensland, New South Wales, and Victoria
(a) Australian Capital Territory Basin ID: 410730
(b) Western Australia Basin ID: 604053
(c) Northern Territory Basin ID: G8200045
(d) Tasmania Basin ID: 318076
Figure 7: Ensemble Quantile-LSTM streamflow predictions for randomly selected basins from Australian Capital Territory, Western Australia, Northern Territory, and Tasmania
(a) Nepean River at Maguires Crossing (Station ID: 212209) in Hawkesbury River region of New South Wales
(b) Barambah Creek at Litzows (Station ID: 136202D) in Burnett river region of Queensland
Figure 8: Flooding events in selected catchments of New South Wales and Queensland predicted using the Ensemble Quantile-LSTM streamflow prediction model
(a) Cullyamurra Water Hole (Station ID: A0030501) in Cooper Creek of South Australia
(b) Darkin River at Pine Plantation (616002) in Swan coast - Avon River region of Western Australia
Figure 9: Flooding events in selected catchments of South Australia and Western Australia predicted using the Ensemble Quantile-LSTM streamflow prediction model
@@ 5 Discussion
In the first set of experiments, we compared the performance of the modelling strategies at the catchment scale using a conventional LSTM model for five-step-ahead (daily) prediction of streamflow rate using climate data. We found that the Individual catchment modelling strategy, where a separate model is used for each gauging station data, provided the best results (Table 3). This is because each catchment has unique characteristics that could be based on weather patterns, climate, and topography. In Table 3, we also observed that the Stacked Ensemble model gives a poor performance, as it cannot beat the individual LSTM models. This could be due to a lack of information flow in combining the time series data with the static model input. We note that static data type is different when compared to time series data and there is a major challenge in machine learning when considering various types of data in data fusion strategies [91] and multiview learning [92]. Comparatively, The Batch-Temporal strategy is essentially the Batch Model without the inclusion of static features. It operates on the same time-series features as the Individual strategy, but with a collective approach, running a single model for all stations in the state rather than individual models for each station. This consolidation aims to optimise computational resources. The Stacked Ensemble strategy obtained predictions from the temporal data using one model and then considered the static data in another model to more accurately capture temporal and static information in an ensemble. The Individual strategy features a specific fit to one catchment at a time and is unaffected by information from other catchments. For example, catchments might vary due to variations in weather patterns, climate, and topography. Overall, these findings are consistent with the work of Leontjeva et al. [26] who similarly compared a variety of strategies i) standalone model using univariate temporal data ii) standalone model using bimodal data (temporal and static) iii) ensemble model using bimodal data. In that study, the standalone model using bimodal data failed to capture information from both sources, while the performance of the ensemble model depended on the distribution of the data.
Generally, it is difficult for a single machine learning model to represent a wide range of data, even if they may be from the same domain, such as in a regionalisation context. Regionalisation would be a challenge for machine learning in general, as a single model cannot be used (fitted to) for several tasks, and in this case, the local catchments can be seen as different tasks as they can have differences in the data distribution dependent on climate/weather/geography. Approaches such as multitask and transfer learning [93, 94] could be considered, where some underlying model knowledge can be re-used and applied to related datasets, i.e. catchments in this case.
In the case of deep learning models using the Individual catchment strategy (Table 4), the Ensemble Quantile-LSTM model provides a mechanism for quantifying uncertainty with close prediction accuracy to conventional LSTM and CNN models. Although designed for image processing purposes, the CNN reproduces streamflow better than conventional LSTM models (Table 4) that were designed for temporal sequences. CNNs may also have advantages for capturing spatial regions (rationalisation) due to their capabilities in image processing, but LSTM models were the best due to their capacity to simulate temporal data. However, Quantile LSTM (q=0.05) in Table 4 provides the best performance accuracy as it has been designed to capture extreme values. The performance of these types of models has been comparable (close), especially for multistep ahead prediction [61]. Furthermore, the Individual catchment strategy does not incorporate the spatial catchment data into the deep learning models, although multivariate data of the local catchment was used. Hence, this could be another reason why LSTM models outperformed the CNN. ED-LSTM and BD-LSTM did not provide the best performance, contrary to past studies for univariate multistep ahead prediction [61]. However, we note that our study featured a multivariate approach with a much larger dataset, hence a direct comparison cannot be done.
The results, reinforce the efficacy of the Ensemble Quantile-LSTM as a robust and adaptable choice for hydrological modelling, specifically for South Australia. This performance highlights the Ensemble Quantile-LSTM model’s resilience and effectiveness in capturing both extreme and regular streamflow which can imply their applicability in predicting flood occurrences. The prediction of extreme values can potentially be corrected with data augmentation with deep learning models. We note that data augmentation and ensemble learning methods have been prominent [25] in class imbalance problems, and we can view the case of extreme value forecasting as a regression case of class imbalance problems.
The predictions essentially give an outlook - up to five days ahead, which has been demonstrated to be useful for flood prediction based on historical flooding events for selected catchments. Although the model performance accuracy deteriorates as the prediction horizon increases in multistep ahead problems [61], it can still provide a forecast of upcoming flooding events and enable better disaster management. The results deteriorate as the gap in the values to predict increases with increasing prediction horizon. In an operational sense, we can envision a model that can be run repeatedly in real-time, providing more and more accurate predictions as more data becomes available in time. We have demonstrated that the streamflow prediction can be used as a means to enable forecasting of floods, as done in past studies [95, 96, 97, 98].
@@
Despite the success of evaluating several machine learning approaches for predicting flood extremes, this study has some obvious limitations. Firstly, the models do not naturally project model uncertainties into the predictions, which is vital for natural disasters caused by extreme events. This can be addressed by Bayesian deep learning [99, 100] and variational deep learning models [101] in future work. Secondly, we have not accounted for topography information in the data to develop a robust spatiotemporal model.
@@
In the case of larger catchments, the steepness and variation in topography can influence rainfall patterns, particularly the within-catchment distribution of the rainfall. In addition, topography and catchment shape can influence the timing between rainfall and flood extremes. Implicitly, the proposed model should be able to capture these variations for the individual catchment strategy based on the data, but it is possible that adding additional information will assist the model in identifying the patterns. We did not incorporate the changes in climate conditions due to the El Niño climate effect, which is a longer scale pattern in the Australian rainfall [102, 103]. Again, the LSTM should be able to discover this from the data, given a long enough time series. However, adding this as a variable might assist the model in identifying the variations in the data. The data is also sparse in space given the locations of the gauging stations (catchments), and this means if we need data for specific regions where no gauging stations are available, generative machine learning models [104], and general circulation models [105] could be used.
@@ 6 Conclusions
In this study, we undertook a critical evaluation of several deep learning models and presented a novel Ensemble Quantile-LSTM model to accurately forecast streamflow extremes over selected Australian catchments using a multi-step-ahead prediction approach. Our investigation revealed that the Individual catchment strategy which features a separate model for a catchment gauging station gives the best performance for extreme value forecasting. Therefore, we found that a tailored approach, where each catchment is paired with a model dedicated to its distinct gauging station, significantly elevated the accuracy of extreme value forecasts. It’s important to highlight that this precision in modelling ensures that local variability and unique hydrological characteristics are captured effectively, providing a more nuanced understanding of each catchment’s behaviour in the face of extreme weather events.
The investigation unveiled the potential of our Ensemble Quantile-LSTM model that integrates quantile regression to surpass all baseline machine learning models for streamflow prediction. This model emerges as a superior tool not only due to its performance but also because it represents a leap towards a more reliable prediction by embracing the inherent uncertainties in hydrological forecasting and addressing the problem of extreme value forecasting. Looking ahead, we recognise that the path to impeccable forecasts is iterative and progressive. Our research underscores the vast room for improvement that can be addressed by extending our dataset and refining our model, particularly in the area of uncertainty quantification. The aspiration is to reach a point where our model can offer robust predictions with quantified confidence levels, transforming how stakeholders prepare for and respond to the challenges posed by extreme rainfall and associated streamflow conditions such as floods. Our Ensemble Quantile framework can be improved further for uncertainty quantification by combining it with Bayesian deep learning.
@@
    Developers: Siddharth Khedkar, Arpit Kapoor and Rohitash Chandra
@@
A. Kapoor, S. Khedkar and J. Ng contributed to coding, writing and experiments. W. Vervoort contributed to analysis, editing and supervision. R. Chandra contributed to conceptualisation, writing, code analysis, and supervision.